# SPECIAL RELATIVITY

### v9

@@ 11.4 Wigner Rotation and the Trouton-Noble Experiment
11.5 Bibliography and Notes
1. The measurements of the resonance absorption of gamma rays in a Mossbouer rotating disk experiment [43] for the same relative velocity between the source and the observer give a blue shift when the source is at the center, and a red shift when it is at the tip.
2. In Ives and Stilwell experiment [44] the effect which occurred on the ground of the orbital motion of the earth relative to the sun was within the experimental limitations. This situation pertains because the high value of the emitters speed $w\gg v$ involved. Presently we give a general analysis of this type of experiments independently of the experimental accuracy.
3. Many physicists tend to think that the validity of Maxwell’s laws in all inertial frames implies the elimination of the privilege frame. To quote e.g. Dieks [45]: ” As we already mentioned, it is a basic principle of the special theory of relativity that the line element $ds$ supplies all information about the physics of the situation, as described in the given coordinates.”
4. The first experimental confirmation of the time dilation predicted in the circle path geometry with atomic clocks and aircraft was reported by Hafele and Keating in 1971 [46]. A very interesting result have been obtained by GPS team. The first example of the different inertial frames having the same accelerational histories is the global positioning system (GPS). The number of GPS satellites is around 80. The velocity of the GPS satellites with regard to the earth is 4 km/s. The experimental data of the GPS shows that the clocks in the satellites tick off time more slowly (7.1 $\mu$s every day) by the velocity. Asymmetry in clock run appears between the earth and the satellites. There are no asymmetries among the GPS satellites. That is, times are equal in every GPS satellites [47]. We see that the presence of some non zero relative velocity cannot be a cause for the effect of time dilation. In this experiment no effect of earth orbital velocity on the clock dilation was found. The negative result of the GPS experiment can be understood remembering that in this situation the clocks also are moving around the earth (i.e. around the reference clock) along the circular path and no observable effect of orbital velocity remains.
@@ 12 The Principle of Relativity and the Modern Cosmology
12.1 The Relativity Principle of Special Relativity
In the last few chapters we have treated the essential ideas necessary for an understanding of the aberration of light phenomena. We have not had to worry yet about the fact that the source-observer asymmetry associated with the stellar aberration phenomenon at first glance contradicts the principle of relativity.
@@ 12.1 The Relativity Principle of Special Relativity
The principle of relativity which we owe to Poincare, who first coined the term, denied the possibility for an observer partaking in a uniform motion relative to the fixed stars of discovering, by any measurement, such a motion, under the assumption, of course, that one does not look outside.
@@
Now let us to analyze the stellar aberration experiment. Perhaps we should say: ” In the case of stellar aberration, the peculiarity of the aberration of light measurements is that the accelerated (earth-based) observer looks outside to the stars.” But no! The motion of the stars with respect to the earth is never followed by any aberration. The aberration shift (as inferred from astronomical observations) exists even in the case when star moves with the same velocity as the earth. We cannot remain with the framework of ”looking outside to the star” when considering the earth-based observations of the stellar aberration.
In the Section 8.3 we discussed a simple scaling model for the stellar aberration. We obtained a condition for optical similarity between the aberration of light from a distant star and from the earth-based incoherent source. We find that without looking outside, it is possible to determine the speed of the earth around the sun by means of aberration of light measurements. No one has ever done such experiment with earth-based incoherent source, but we know would happen from the astronomical observations.
It should be note that all well known methods to test the special relativity are second order measurements . The cardinal example is given by the Michelson-Morley experiment. A close examination of all second order experiments inside an inertial frame shows that phenomena appear to be independent of the uniform motion relative to the fixed stars. From the series of negative results like that obtained by Michelson and Morley it seems reasonable to suppose that it is not accidental that various optical experiments trying to determine orbital velocity remained unsuccessful. And we should accept the view that there is exist a general law of of nature which prevents us to determine orbital velocity by laboratory experiments. Nevertheless, in this case there does not appear to be any such profound implication. This is specific feature of the optical experiments of second order. All these methods are based on phenomenon of interference. There is a fact governing all interference phenomena. The law is called the relativistic invariance of phase (i.e. it is simply number). It is not change under any reference frame transformations. It is not possible to detect the orbital velocity using second order interference experiments. This is consequence of the geometry of space-time. We conclude the following: The usual formulation of the theory of relativity deals with formulated above principle of irrelevancy of velocity, but it should be understood only in a limiting sense when we are dealing with second order optical measurements.
@@
There is no conflict between the fundamental structure of special theory of relativity on the one hand, and the aberration of light phenomena. First, from what is said in the previous chapters it should be clear that special relativity does not require the formulated above principle of irrelevancy of velocity. Principle of special relativity applies to physical laws, not to physical facts. Not only in Einstein writings, but in every textbook on special theory of relativity can be found the formulation of principle of special relativity as follows: the laws of physics are the same in all inertial frames (1).
The special principle of relativity says that the same laws should hold in all inertial frames. We interpret this to mean that the laws can be expressed by equations that have the same form in all inertial frames. According to special principle of relativity all inertial frames are equivalent with regard to physical laws, not with regard to physical facts. By a physical fact in this context we mean properties of existing objects - in our case of interest the properties of incoherent source. By mean a physical law we mean for example electrodynamics equations. Principle of special relativity is irrelevancy of velocity with regard to physical laws, not with regard to anything.
We point that the essence of the special theory of relativity consists in the following postulate: all physical process proceed in four-dimentional space-time, the geometry of which is pseudo-Euclidean. But give pseudo-Euclidean geometry of space-time does not dictate the irrelevancy of velocity with regard to anything. From a mathematical viewpoint, the argument looks something like the following. The inertial frames are characterized by motion at constant velocity relative to the fixed stars and different inertial frames are related by Lorentz boosts. Relativistic symmetry is usually identified with Lorentz symmetry, i.e. with symmetry under Lorentz boosts. At the same time it is common knowledge that Lorentz boosts alone do not make up a group. Lorentz boosts do not form a group due presence of the Wigner rotation. In general, two successive Lorentz boosts are not equivalent to boost. The product of two Lorentz boosts in different directions is equal to the product of a pure boost and a spatial rotation, the Wigner rotation. We must conclude that when we are dealing with transformations and set of these transformations does not form a group we have no symmetry between inertial frames. There is a real difference between the accelerated inertial frame and the inertial frame without accelerational history. It is comes out quite natural that this difference is closely associated with the Wigner rotation phenomenon.
@@
The existence in nature of something corresponding to the concept of absolute acceleration entails the existence in nature of something corresponding to the concept of absolute velocity. If we look at the situation carefully we see that the time dilation effect problem in the accelerated frame demonstrate the essential asymmetry between accelerated and initial ( i.e. without acceleration history) inertial frame. The proper time of any particle moving arbitrary in the initial inertial frame is always slower than the physical time in this initial inertial frame. It is a fundamental property of the pseudo-Euclidean space-time. There is no relativity of the uniform motions and we can determine the particle absolute velocity in the initial inertial frame. The formalism of classical physics is associated with absolute space structure. This structure must be associated with some ”substrance”, the ether, which incidentally provides an absolute rest frame. So is space-time the new ether? It does have similar properties. It clearly does provide an ”absolute rest reference frame”. The main difference between the initial inertial frame and the absolute space of Newton is that special relativity taught us to think in terms of unified space-time model, with no a priori splitting of space from time. The initial inertial frame is not physically identical to the absolute space. It is a concept of an absolute space-time.
@@
The relativity principle has a long history in physics. Newtonian mechanics is invariant under Galilean transformations. However, from the mathematical viewpoint, the relativistic Lorentz transformation is qualitatively different from the non-relativistic Galilean one. The set of Galileo boosts form a group. So we see that principle of Galilean relativity states that all inertial frames are equivalent. Equivalence between two frames means that none of them preferred to the other. Galileo’s version of relativity principle was incorrectly extended to the whole of physics. The relativity principle does not hold for the whole range of validity of the Lorentz covariant physical laws and the Lorentz covariance is not a fundamental symmetry of physics (because the accelerated inertial frame is connected with the initial inertial frame by the Galilean transformation, but not the Lorentz one).
12.2 Aberration and the Modern Cosmology
@@
Now we want to discuss a serious trouble - the failure of the aberration of light theory presented above. There are difficulties associated with the ideas of modern cosmology. The difficulty we speak of is associated with the concept of source-observer asymmetry, when applied to the motion relative to the cosmic microwave background (CMB). One may say that perhaps there is no use worrying about these difficulties since there are so many things about universe that we still don’t understand. It is obvious that we live in a universe ruled by the unknown. Indeed, dark energy and dark matter together contribute to almost 96 percents of the total energy-mass of the universe and we have no knowledge of what dark matter (or dark energy) is. It is important to realize that when we try to talk about working assumptions in modern cosmology, they are approximate and will change. Therefore what we discuss in this section will not be accurate in a certain sense.
@@
We cannot go into details of the modern cosmology principles at this time. We shall assume that they are there, and go on to describe what some of the consequences are. According to the cosmological principle, the universe should appear isotropic, without any preferred direction, to a comoving observer, having no peculiar motion relative to the cosmic fluid of the expanding universe. However, a peculiar motion of such an observer might introduce a dipole anisotropy in the observed properties of a class of objects and which in turn, might be exploited to infer the peculiar velocity of the observer. For instance the CMB radiation shows an anisotropy distribution. The CMB dipole is almost ubiquitously assumed to be of kinematical origin, i.e. due to relative motion. It is widely believed that the dipole anisotropy is produced by the Doppler effect due to the relative motion between the earth, i.e. an observer, and the frame where cosmic microwave background looks nearly isotropic. By subtracting the dipole, the CMB is defined as the rest frame for the universe.
@@
The observed dipole indicates that the solar system is moving at 370 km/s relative to the observed universe in the direction of galactic longitude $l=264^{o}$ and latitude $b=48^{O}$. This is quite far from the galactic rotation direction 250 km/s towards $l=90^{o}$ and $b=0$. The motion relative to the cosmic microwave background results from the sum of many components of velocity due to gravitational attraction of various mass concentrations. The existence of clusters and super clusters of galaxies and our motion is a natural consequence of the large scale organization of matter. The peculiar velocity consists of the five vector contributions: the motion of the earth around the sun (30 km/s), the hypothetical circular motion around the our galaxy (250 km/s), the motion of our galaxy in the local group, and the motion of the local group with respect to the cosmic microwave background. The origin of the velocity of the local group is still uncertain and has been under discussion over past two decades. This peculiar velocity is believed to be generated by the spatial inhomogeneties of mass (mainly dark matter) distribution in nearly large scale structures. The velocity of the local group with respect to the cosmic microwave background is estimated to be 500 km/s.
@@
Now let us see how aberration of light various with the speed relative to the rest frame of the universe. Consider first the textbook explanation. It is very easy to understand how the aberration of light effect comes about from the point of view of the conventional aberration of light theory. The existence of an aberration of the line of sight by motion can be recognized by considering an observer in a car driving through rainstorm: raindrops falling vertically appear to be obliquely. An important application of the Galilean velocity transformation law is provided by stellar aberration, the change in the apparent direction of a star caused by the earth’s motion around the sun. The apparent positions of all fixed stars are thus always a little displaced in the direction of the earth’s motion at that moment, and hence describe a small elliptical figure during the annual revolution of the earth around the sun. What about the motion relative to the CMB rest frame? According to textbooks, if the earth motion were uniform, the aberration effect would be undetectable since the ”true” direction of the star is unknown. Indeed, who can say where a given star should be? On the other hand, the true direction of an earth-based source is known. But, according to conventional theory, the aberration of light phenomenon does not exist in an aberration of light experiments using an earth-based light source.
@@
We now in position to understand the nature of our difficulty with the motion relative to the CMB. It is believed that the dipole anisotropy is produced by the Doppler effect due to acceleration (during the billions of years) of the solar system with respect to the rest frame of the universe. This acceleration is believed to be generated by the spatial inhomogenetics of (Dark matter) mass distribution in nearly large scale structures. A correct solution of the aberration problem in the earth-based frame requires the use of metric tensor even in first-order experiments since the crossed term in the Langevin metric plays a fundamental role in the non-inertial kinematics of a light ( or relativistic particle ) beam produced by the earth-based source. According to the relativistic theory of aberration presented in this book, the proper rotation of the earth on its axis should produce a corresponding shift of the image. The aberration shift will depend also upon the value of $v_{\perp}^{CMB}$, the component of the solar system velocity (relative to the CMB) perpendicular to the earth rotation axis. The rotation of the earth produces aberration in an amount larger enough to be taken into account in precise observation work using electron microscope as an earth-based particle source. Experimental results show (see the Chapter 8) that the image shift is quite close to the theoretical prediction for the (30 km/s) orbital velocity and clearly indicate that the signal associated with motion (370 km/s) relative to the CMB does not exist. The simplest explanation is that the CMB dipole might be of non-kinematical origin.
@@
More recent observations of astronomers also cast doubt on the CMB dipole, being the ultimate representatives of the solar peculiar velocity. In recent years observations have emerged hinting at an anisotropic universe. The discovery of the preferred direction in the universe was serendipitous. Until very recently the velocity of the solar system in the rest frame of universe is inferred from CMB temperature dipole anisotropy. Obviously, an independent measurements of this velocity is needed to fully establish the kinematical origin of the dipole. Another such quantity that could be employed to look for departures from isotropy is the angular distribution of distant radio sources in the sky. This could provide an independent check on the interpretation of CMB dipole. The radio data clearly indicate that significantly larger dipole exist in the rest frame of the radio galaxies. While the velocity of the solar system inferred from the CMB temperature dipole anisotropy is 370 km/s, the radio dipole measurements finds the speed of motion to be around 1000 km/s (i.e. to be around three times larger that that of CMB). From the Hubble diagram of quasars, motion of the solar system is derived, which out to be the largest value ever found, 8000 km/s (2).
@@
On the other hand, a common direction for all these dipoles, determined from completely independent surveys by different groups employing different techniques, indicate that these dipoles are not resulting from some systematics in the observations or the data analysis, but could instead suggest an inherent anisotropy. This is totally unexpected in a standard model of universe. We have a preferred direction, aligned with the CMB dipole, in the universe. That is, going to the CMB rest frame, we see an anisotropic background. There is a difficulty with such some sort of an ”axis” of the universe which, in turn, would be against the cosmological principle. Three independent dipole vectors pointing along the same particular direction could imply an anisotropic universe, violating the cosmological principle, a cornerstone of the modern cosmology(2).
@@
Let us be conservative and say that there are two kinds of the earth velocity that the total velocity with respect to the initial (i.e. privilege ) inertial frame could be the sum of the orbital velocity and the velocity of the sun. The velocity of the sun consists the two contributions: the circular motion around our galaxy and the motion of our galaxy. In earth-based experiments where we measure the earth velocity with respect to the privilege frame by seeing aberration of light (or particles), we are measuring the recorded (in the accelerational history) velocity. So the velocity with respect to the privilege frame consists of the two contributions: a recorded motion plus an unrecorded motion. We know that there is definitely a recorded (orbital) motion, and we have a formula for it. It is therefore impossible to get all the earth velocity to be record in the accelerational history in the way we hoped. It is not a legal theory if we have nothing but modern cosmology. Something else has to be added.
@@
We would like to think a little more about why the earth is moving with respect to the initial inertial frame. There must have been a force pushing on the earth in order to get it going. So it may help our understanding if we look a little more closely at where the forces come from. We must say immediately that orbital motion is result of non-gravitational forces. It is well known that a full 98 percent of all the angular momentum in the solar system is concentrated in the planets, yet a staggering 99.8 percent of all the mass in our solar system is in our sun. Perhaps the first scientifically respectable theory of the origin of the solar system was given by Hoyle (1960). He invoked the action of magnetic field to transfer angular momentum from the central body, the sun, to the ejected matter which eventually formed the planets. In contrast, the unrecorded motion around our galactic and the motion of our galactic is result of the action of gravity.
@@
We started by talking about the gravitational interaction of the earth with the sun. We believe the theory of gravity is so much that we allow it to tell us about the force of one galactic to another. Perhaps we are making to great an extrapolation of our limiting knowledge of gravity to galactic scales. Perhaps the entire difficulty is that a modification of gravity may be responsible for unrecorded velocities. We may someday find out one things we don’t understand today, for example dark matter, can, in fact be explained as modification of the gravitation theory. Today it doesn’t seem likely, but no one can say for sure. There are so many things about universe that we don’t understand.
12.3 Bibliography and Notes
1. By ”the same” is means form-invariance: the mathematical relations between the physical quantities remain identical after a coordinate transformation has been performed on these quantities. In special theory of relativity the (form-invariant) transformations between inertial frames are Lorentz-transformations [48].
2. Recent observations of the cosmic dipole anisotropies is revolutionized our understanding of the universe. Attempts to recover the CMB dipole from counts of later universe sources such as radio galaxies and quasars, which are assumed to be in the CMB frame, largely agree that CMB dipole direction is recovered, but not the magnitude. The various dipoles, included CMB dipole, all pointing along the same direction, suggest a preferred direction in the universe, raising thereby uncomfortable questions about the cosmological principle, the basis of the standard model in modern cosmology [49, 50, 51]. .

### v10

@@ 15 The Principle of Relativity and the Modern Cosmology
Above we have treated the essential ideas necessary for an understanding of the aberration of light phenomena. At first glance, the source-observer asymmetry associated with the stellar aberration phenomenon contradicts the principle of relativity. This principle which we owe to Poincare, who first coined the term, denied the possibility for an observer partaking in a uniform motion relative to the fixed stars of discovering, by any measurement, such a motion, under the assumption, of course, that one does not look outside.
@@
Now let us analyze the stellar aberration experiment. Perhaps we should say: ” In the case of stellar aberration, the peculiarity of the aberration of light measurements is that the accelerated (earth-based) observer looks outside to the stars.” However, the motion of the stars with respect to the earth is never followed by any aberration. The aberration shift (as inferred from astronomical observations) exists even in the case when a star moves with the same velocity as the earth. We cannot remain within the framework of ”looking outside to the star” when considering the earth-based observations of the stellar aberration. In Chapter 10 we discussed a simple scaling model for stellar aberration. We obtained a condition for optical similarity between the aberration of light from a distant star and from the earth-based incoherent source. We find that without looking outside, it is possible to determine the speed of the earth around the sun using aberration of light measurements.
@@
There is no conflict between the fundamental structure of the special theory of relativity on the one hand, and the aberration of light phenomena. First, from what is said in the previous chapters it should be clear that special relativity does not require the formulated above principle of irrelevancy of velocity. The principle of special relativity applies to physical laws, not to physical facts. The special principle of relativity says that the same laws should hold in all inertial frames. We interpret this to mean that the laws can be expressed by equations that have the same form in all inertial frames  [^57] In the special theory of relativity the form-invariant transformations between inertial frames are Lorentz transformations [57]. For a general discussion about the Principle of Relativity we suggest reading the paper [58]. . According to the special principle of relativity, all inertial frames are equivalent with regard to physical laws, not with regard to physical facts.
@@
The existence in nature of the concept of absolute acceleration entails the existence in nature of the concept of absolute velocity. If we look at the situation carefully we see that the time dilation effect problem in the accelerated frame demonstrates the essential asymmetry between accelerated and initial ( i.e. without acceleration history) inertial frame. The proper time of any particle moving arbitrarily in the initial inertial frame is always slower than the physical time in this initial inertial frame. It is a fundamental property of the pseudo-Euclidean space-time. There is no relativity of the uniform motions and we can determine the particle’s absolute velocity in the initial inertial frame. The formalism of classical physics is associated with absolute space structure. This structure must be associated with some ”substance”, the ether, which incidentally provides an absolute rest frame. So is space-time the new ether? It does have similar properties. It clearly does provide an ”absolute rest reference frame”. The main difference between the initial inertial frame and the absolute space of Newton is that special relativity taught us to think in terms of a unified space-time model, with no a priori splitting of space from time. The initial inertial frame is not physically identical to the absolute space. It is a concept of an absolute space-time.
@@
The relativity principle has a long history in physics. Newtonian mechanics is invariant under Galilean transformations. However, from the mathematical viewpoint, the Lorentz transformation is qualitatively different from the Galilean one. It is common knowledge that Lorentz boosts alone do not make up a group. The set of Galileo boosts form a group. So we see that the principle of Galilean relativity states that all inertial frames are equivalent. Equivalence between two frames means that none of them is preferred to the other. The reason for this (Galilean group) symmetry is simply that equations of motion do not explicitly depend on velocity. The internal dynamics of the system do not change under the Galilean boost. Galileo’s version of the relativity principle was incorrectly extended to the whole of physics.
The relativity principle does not hold for the whole range of validity of the Lorentz covariant physical laws and the Lorentz covariance is not a fundamental symmetry of physics. In fact, since the metric tensor of space-time must be a continuous quantity, the accelerated inertial frame is connected with the initial inertial frame by the Galilean transformation, but not the Lorentz one. The Galilean transformations do not preserve the form-invariance of Maxwell’s equations under a change of inertial frame. The reason for this absence of Galileo’s relativity is simply that electrodynamics equations explicitly depend on velocity. Therefore, according to special relativity, there is a real difference between the accelerated (with respect to the fixed stars) inertial frame and the inertial frame without accelerational history. It should be noted that this difference is closely related to the inertial forces. This is the basic content of the equivalence principle. The relation between inertia and gravity is a subject that needs to be understood more deeply.
@@
Now we want to discuss a serious difficulty associated with the ideas of modern cosmology. The difficulty we speak of is associated with the concept of source-observer asymmetry when applied to the motion relative to the cosmic microwave background (CMB). One may say that perhaps there is no use worrying about these difficulties since there are so many things about the universe that we still don’t understand. We live in a universe ruled by the unknown. Indeed, dark energy and dark matter together contribute to almost 96 percent of the total energy-mass of the universe and we have no knowledge of what dark matter (or dark energy) is. It is important to realize that when we try to talk about working assumptions in modern cosmology, they are approximate and will change.
@@
We cannot go into details of the modern cosmology principles at this time. We shall assume that they are there, and go on to describe what some of the consequences are. According to the cosmological principle, the universe should appear isotropic, without any preferred direction, to a comoving observer, having no peculiar motion relative to the cosmic fluid of the expanding universe. However, a peculiar motion of such an observer might introduce a dipole anisotropy in the observed properties of a class of objects which in turn, might be exploited to infer the peculiar velocity of the observer. For instance, the CMB radiation shows an anisotropy distribution. The CMB dipole is almost ubiquitously assumed to be of kinematical origin, i.e. due to relative motion. It is widely believed that the dipole anisotropy is produced by the Doppler effect due to the relative motion between the earth, i.e. an observer, and the frame where the cosmic microwave background looks nearly isotropic. By subtracting the dipole, the CMB is defined as the rest frame for the universe.
@@
The observed dipole indicates that the solar system is moving at 370 km/s relative to the observed universe in the direction of galactic longitude $l=264^{o}$ and latitude $b=48^{O}$. This is quite far from the galactic rotation direction of 250 km/s towards $l=90^{o}$ and $b=0$. The motion relative to the cosmic microwave background results from the sum of many components of velocity due to the gravitational attraction of various mass concentrations. The existence of clusters and super clusters of galaxies and our motion is a natural consequence of the large-scale organization of matter. The peculiar velocity consists of the five vector contributions: the motion of the earth around the sun (30 km/s), the hypothetical circular motion around our galaxy (250 km/s), the motion of our galaxy in the local group, and the motion of the local group with respect to the cosmic microwave background. The origin of the velocity of the local group is still uncertain and has been under discussion over the past two decades. This peculiar velocity is believed to be generated by the spatial inhomogeneities of mass (mainly dark matter) distribution in nearly large-scale structures. The velocity of the local group with respect to the cosmic microwave background is estimated to be 500 km/s.
@@
Now let us see how aberration of light varies with the speed relative to the rest frame of the universe. Consider first the textbook explanation. It is very easy to understand how the aberration of light effect comes about from the point of view of the conventional aberration of light theory. The existence of an aberration of the line of sight by motion can be recognized by considering an observer in a car driving through a rainstorm: raindrops falling vertically appear to be oblique. An important application of the Galilean velocity transformation law is provided by stellar aberration, the change in the apparent direction of a star caused by the earth’s motion around the sun. The apparent positions of all fixed stars are thus always a little displaced in the direction of the earth’s motion at that moment, and hence describe a small elliptical figure during the annual revolution of the earth around the sun. What about the motion relative to the CMB rest frame? According to textbooks, if the earth’s motion were uniform, the aberration effect would be undetectable since the ”true” direction of the star is unknown. Indeed, who can say where a given star should be? On the other hand, the true direction of an earth-based source is known. But, according to conventional theory, the aberration of light phenomenon does not exist in an aberration of light experiments using an earth-based light source.
@@
We are now in a position to understand the nature of our difficulty with the motion relative to the CMB. It is believed that the dipole anisotropy is produced by the Doppler effect due to acceleration (during the billions of years) of the solar system with respect to the rest frame of the universe. This acceleration is believed to be generated by the spatial inhomogeneities of (Dark matter) mass distribution in nearly large-scale structures. A correct solution of the aberration problem in the earth-based frame requires the use of metric tensor even in first-order experiments since the crossed term in the Langevin metric plays a fundamental role in the non-inertial kinematics of a light ( or relativistic particle ) beam produced by the earth-based source. According to the relativistic theory of aberration presented in this article, the proper rotation of the earth on its axis should produce a corresponding shift of the image. The aberration shift will depend also upon the value of $v_{\perp}^{CMB}$, the component of the solar system velocity (relative to the CMB) perpendicular to the earth’s rotation axis. The rotation of the earth produces aberration in an amount larger enough to be taken into account in precise observation work using an electron microscope as an earth-based particle source. Experimental results show (see Chapter 10) that the image shift is quite close to the theoretical prediction for the (30 km/s) orbital velocity and clearly indicate that the signal associated with motion (370 km/s) relative to the CMB does not exist. The simplest explanation is that the CMB dipole might be of non-kinematical origin.
@@
More recent observations by astronomers also cast doubt on the CMB dipole, being the ultimate representative of the solar peculiar velocity. In recent years observations have emerged hinting at an anisotropic universe. The discovery of the preferred direction in the universe was serendipitous. Until very recently the velocity of the solar system in the rest frame of the universe is inferred from CMB temperature dipole anisotropy. Obviously, an independent measurement of this velocity is needed to fully establish the kinematical origin of the dipole. Another such quantity that could be employed to look for departures from isotropy is the angular distribution of distant radio sources in the sky. This could provide an independent check on the interpretation of CMB dipole. The radio data clearly indicate that significantly larger dipoles exist in the rest frame of the radio galaxies. While the velocity of the solar system inferred from the CMB temperature dipole anisotropy is 370 km/s, the radio dipole measurements find the speed of motion to be around 1000 km/s (i.e. to be around three times larger than that of CMB). From the Hubble diagram of quasars, the motion of the solar system is derived, which is out to be the largest value ever found, $8\cdot 10^{3}$ km/s  [^58] Recent observations of the cosmic dipole anisotropies is revolutionized our understanding of the universe. The various dipoles, including CMB dipole, all pointing along the same direction, suggest a preferred direction in the universe, raising thereby uncomfortable questions about the cosmological principle, the basis of the standard model in modern cosmology [59, 60, 61]..
@@
On the other hand, a common direction for all these dipoles, determined from completely independent surveys by different groups employing different techniques, indicates that these dipoles are not resulting from some systematics in the observations or the data analysis, but could instead suggest an inherent anisotropy. This is totally unexpected in a standard model of the universe. We have a preferred direction, aligned with the CMB dipole, in the universe. That is, going to the CMB rest frame, we see an anisotropic background. There is a difficulty with such some sort of an ”axis” of the universe which, in turn, would be against the cosmological principle. Three independent dipole vectors pointing along the same particular direction could imply an anisotropic universe, violating the cosmological principle, a cornerstone of modern cosmology.
@@
Let us be conservative and say that there are two kinds of the earth velocity. The total velocity with respect to the initial (i.e. privilege ) inertial frame could be the sum of the orbital velocity and the velocity of the sun. The velocity of the sun consists the two contributions: the circular motion around our galaxy and the motion of our galaxy. In earth-based experiments where we measure the earth’s velocity with respect to the privileged frame by seeing the aberration of light (or particles), we are measuring the recorded (in the accelerational history) velocity. So the velocity with respect to the privilege frame consists of two contributions: a recorded motion plus an unrecorded motion. We know that there is definitely a recorded (orbital) motion. It is therefore impossible to get all the earth’s velocity to be recorded in the accelerational history in the way we hoped. Clearly something else has to be added.
@@
We would like to think a little more about why the earth is moving with respect to the initial inertial frame. There must have been a force pushing on the earth in order to get it going. So it may help our understanding if we look a little more closely at where the forces come from. We must say immediately that orbital motion is the result of non-gravitational forces. It is well known that a full 98 percent of all the angular momentum in the solar system is concentrated in the planets, yet a staggering 99.8 percent of all the mass in our solar system is in our sun. Perhaps the first scientifically respectable theory of the origin of the solar system was given by Hoyle (1960) [62]. He invoked the action of the magnetic field to transfer angular momentum from the central body, the sun, to the ejected matter which eventually formed the planets. In contrast, the unrecorded motion around our galactic and the motion of our galactic is the result of the action of gravity.
@@
We started by talking about the gravitational interaction of the earth with the sun. We believe the theory of gravity is so much that we allow it to tell us about the force of one galactic to another. Perhaps we are making too great an extrapolation of our limiting knowledge of gravity to galactic scales. Perhaps the entire difficulty is that a modification of gravity may be responsible for unrecorded velocities. There are so many things about the universe that we don’t understand.